# TACO: Topics in Algorithmic COde generation dataset

### 22 Dec 2023 v1

@@ 3.1 Data collection phase
To guarantee the dataset’s integrity and the comprehensiveness of its tagging metadata, we devised bespoke HTML parsers tailored to each of the four major programming contest platforms: CodeChef, CodeForces, HackerRank, and GeeksforGeeks. These parsers are engineered to accommodate the diverse syntactic layouts present in the problem texts of their source websites. Throughout the parser development phase, we confirmed the accuracy of each specialized parser via extensive manual sampling. On exceptional occasions, we also employed OCR tool APIs (e.g., SimplePix) to facilitate the conversion of svg-formatted information into LaTeX format. For a detailed discussion of validation procedures and data scraping techniques used for these platforms. Additionally, with regard to the algorithmic labeling designed to augment the APPS and CodeContest datasets, our HTML parser similarly crawls algorithmic skill labels from sources such as LeetCode, Kattis, CodeWars, and CodeForces.
@@ 3.2.1 Correctness processing
  • Unit Test Validation: While the code harvested from programming contest platforms was ostensibly marked as “correct” by those sites, our examination revealed instances where some code either was incorrect or failed to meet predefined criteria. To guarantee that the code within our dataset met a robust standard of correctness, we executed each solution against a comprehensive set of predefined unit tests, subsequently omitting any that failed to pass this rigorous evaluation.
@@ 3.2.2 Functional processing
To meet our research objectives, we employed the aforementioned methodology and executed meticulous data processing on the multiple components of the TACO dataset, as delineated in Huangtao: Table 3. Specifically, we undertook a rigorous data-cleaning process for the newly acquired problems from CodeChef, CodeForces, HackerRank, and GeeksforGeeks. Likewise, we conducted an exhaustive cleaning procedure on problems obtained from open-source datasets—namely APPS, CodeContest, and Description2code—with the objective of eliminating any invalid information from the solutions, thereby enhancing the accuracy and reliability of the TACO dataset.
@@ 4 Comparisons with Code Datasets
Regarding model validation, APPS principally employs GPT-2 and GPT-Neo as benchmark models, thereby neglecting an evaluation of the fine-tuning capabilities of contemporary code models, specifically in the context of programming challenges. In contrast, our TACO dataset comprises a suite of benchmarks generated through the fine-tuning of state-of-the-art models; notably, the most extensive among these employs the StarCoder model, which boasts 15 billion parameters Huangtao: (refer to Section x.x for further details).

### 25 Dec 2023 v2

@@ 3.1 Data collection phase
To guarantee the dataset’s integrity and the comprehensiveness of its tagging metadata, we devised bespoke HTML parsers tailored to each of the four major programming contest platforms: CodeChef, CodeForces, HackerRank, and GeeksforGeeks. These parsers are engineered to accommodate the diverse syntactic layouts present in the problem texts of their source websites. Throughout the parser development phase, we confirmed the accuracy of each specialized parser via extensive manual sampling. On exceptional occasions, we also employed OCR tool APIs (e.g., SimplePix) to facilitate the conversion of svg-formatted information into LaTeX format. For a detailed discussion of validation procedures and data scraping techniques used for these platforms, refer to appendix A. Additionally, with regard to the algorithmic labeling designed to augment the APPS and CodeContest datasets, our HTML parser similarly crawls algorithmic skill labels from sources such as LeetCode, Kattis, CodeWars, and CodeForces.
@@ 3.2.1 Correctness processing
  • Unit Test Validation: While the code harvested from programming contest platforms was ostensibly marked as “correct” by those sites, our examination revealed instances where some code either was incorrect or failed to meet predefined criteria, such as time or space complexity, time or space limitations, or adherence to coding standards. To guarantee that the code within our dataset met a robust standard of correctness, we executed each solution against a comprehensive set of predefined unit tests that assessed these specific criteria, subsequently omitting any that failed to pass this rigorous evaluation.
@@ 3.2.2 Functional processing
To meet our research objectives, we employed the aforementioned methodology and executed meticulous data processing on the multiple components of the TACO dataset, as delineated in Table 2. Specifically, we undertook a rigorous data-cleaning process for the newly acquired problems from CodeChef, CodeForces, HackerRank, and GeeksforGeeks. Likewise, we conducted an exhaustive cleaning procedure on problems obtained from open-source datasets—namely APPS, CodeContest, and Description2code—with the objective of eliminating any invalid information from the solutions, thereby enhancing the accuracy and reliability of the TACO dataset.
@@ 4 Comparisons with Code Datasets
Regarding model validation, APPS principally employs GPT-2 and GPT-Neo as benchmark models, thereby neglecting an evaluation of the fine-tuning capabilities of contemporary code models, specifically in the context of programming challenges. In contrast, our TACO dataset comprises a suite of benchmarks generated through the fine-tuning of state-of-the-art models; notably, the most extensive among these employs the StarCoder model, which boasts 15 billion parameters (refer to Section 5.2 for further details).